OpenStudy (anonymous):

use Gauss's approach to find the follwoing sums (do not use formulas). a... 1+2+3+4+...+999 b... 1+3+5+7+...+997 the sum of sequance a is... the sum of sequance b is...

OpenStudy (anonymous):

3+997 = 1000

OpenStudy (anonymous):

i forgot how im suppose to solve this though

OpenStudy (anonymous):

notice that you can sum up pairs of numbers such that their sum is 1000. ( 1 + 999 ) = 1000, ( 2 + 998 ) = 1000, ( 3 + 997 ) = 1000 count hot many time you can go this and multiply by 1000

OpenStudy (anonymous):

but i dont know wat im supose to write for the answer im still lost lol

OpenStudy (anonymous):

\[1\;\; \;\;\;+\;\; \;\;\;2\;\; \;\;\;+\;\; \;\;\;3\;\; \;\;\;+\;\; \;\;\;...........+996+997+998+999\] \[999\;\; \;+\;\; \;998\;\; \;+\;\; \;997\;\; \;+\;\; \;\;\;.........+\;\; \;4\;\; \;+\;\; \;3\;\; \;+\;\; \;2\;\; \;+\;\; \;1\] \[1000\;\;+\;\;1000\;\;+\;\;1000\;\;+\;\;.........+\;\;1000\;\;+\;\;1000\;\;+\;\;1000\;\;+\;\;1000\] So, you've got 999 numbers, and all of them are 1000. So \[\sum_{k=0}^{999} \;n= \frac{999*1000}{2}\] Since you counted the sum twice, look at it carefully, I divided bu two.

OpenStudy (anonymous):

499500??

OpenStudy (anonymous):

yes